# Code Digital Twin: Empowering LLMs with Tacit Knowledge for Complex Software Maintenance

### v1

@@ 1. Introduction
Recent advancements in large language models (LLMs) have demonstrated their impressive capabilities in software engineering (SE) tasks, such as code completion, bug detection, and software documentation. These models excel at understanding and generating code, streamlining development cycles, and assisting developers with routine tasks. Leveraging vast amounts of public domain data, LLMs have notably improved productivity and code quality in simpler, well-defined software systems.
However, as software systems increase in complexity, LLMs face significant challenges. While they are effective in tackling individual tasks, their ability to support the maintenance and evolution of large, intricate systems is limited. LLMs struggle to capture the tacit knowledge underlying complex software—such as high-level functionalities and design decisions—which is often embedded in informal documentation or resides solely in developers’ minds. Moreover, LLMs lack a holistic, system-wide understanding that is necessary to manage the intricate relationships and continuous changes inherent in evolving software.
@@
To address these challenges, we propose the concept and framework of Code Digital Twin, a conceptual representation of tacit knowledge that encapsulates the concepts, functionalities, and design rationales behind code elements, co-evolving with the software. The framework structures tacit knowledge into two key categories: the artifact-oriented backbone, which establishes a foundational structure by linking concepts and functionalities to software artifacts, and the rationale-centric explanations, which enrich this foundation by capturing the decisions that have shaped the system’s design and evolution. The methodology for realizing a code digital twin combines structured and unstructured knowledge representations, integrating knowledge graphs, frames, and textual descriptions to capture domain concepts, functionalities, and rationales. The construction pipeline consists of three key phases: artifact-oriented backbone knowledge extraction, which employs top-down schema-based and schema-free extraction from documentation and bottom-up functionality summarization from source code; rationale-centric explanatory knowledge extraction, where LLMs extract and structure design rationales from unstructured sources like mailing lists and discussions using predefined knowledge frame schemas; and artifact-knowledge reflection construction, which links extracted knowledge across software artifacts and elements to create a cohesive representation. This pipeline supports incremental updates, ensuring that the code digital twin evolves in sync with software changes, maintaining continuous alignment with the latest development state.
@@
By providing integrated knowledge and ensuring continuous alignment with the evolving software, the code digital twin enhances collaboration between LLMs and developers. It enhances LLM capabilities in software maintenance through various modes, including software question answering, context-aware development co-pilot, and autonomous development auto-pilot. This framework presents an innovative approach for systematical accumulation of tacit knowledge in long-term software evolution, enabling more efficient maintenance and adaptation over time.
Furthermore, we outline the challenges and opportunities in the continuous construction of a code digital twin. We discuss the challenges associated with representing, updating, and integrating diverse knowledge elements when applying the construction methodology, ensuring the code digital twin remains relevant and robust as the software evolves. Additionally, we explore the difficulties in leveraging this framework to enhance collaboration between LLMs and human developers, aiming to foster more efficient and adaptive software management.
@@ 2. Motivation
The scale and complexity of software are constantly expanding, while the software developers are also continually changing. It results in an increasing complexity of software development, with related knowledge becoming more intricate. For example, the Linux kernel systems are still undergoing rapid and active development (Linux, 2025), which had approximately 243,000 authors contributing more than 1,060,172 commits in 2021; additionally, its ecosystem is also evolving in a complex manner, involving developers from various countries and organizations. Furthermore, software development knowledge is continuously being lost, as the quality of documentation is often suboptimal, failing to provide and accumulate reliable sources of knowledge. As a result, although LLMs have been extensively trained upon public corpora and have shown promising results for some software development tasks, they struggle to support the development and maintenance of complex software.
Challenges in complex maintenance tasks. Software maintenance is a critical but challenging phase in development lifecycle. On the one hand, maintaining existing code involves difficulties such as pinpointing issues and determining the scope of their impact; on the other hand, new code must be integrated into the complex software architecture and existing code contexts with considering its potential effects. Based on the recent evaluation from OpenAI (Samuel et al., 2025), state-of-the-art LLMs exhibit limited effectiveness in addressing the challenging real-world issue resolution tasks.
Challenges in capturing tacit knowledge. In software requirements and design, a substantial amount of tacit knowledge remains difficult to capture. The success of LLMs mainly stems from their ability to learn from public textual corpora, including code, documentation, or books. However, much of the knowledge related to software requirements and design is not explicitly or well documented. For example, it often resides in the minds of developers, on whiteboards, or in discussions held during project meetings. As a result, existing LLMs have limited capabilities of capturing such important but tacit knowledge. Moreover, such tacit knowledge is difficult and unsuitable for direct injection into LLMs through fine-tuning due to its domain-specific, continuously evolving nature and the needs for human expertise in inspecting and supplementing such knowledge.
Challenges in handling the internal and external complexity of software. Software exhibits both internal and external complexity. On the one hand, design decisions often span over a broad scope, involving intricate interdependencies between various parts of the code repository; on the other hand, designing software also requires consideration of external system factors. However, current LLMs lack the capabilities of higher-level “abstract” thinking and a system-level perspective to handle these complexities.
In summary, LLMs are insufficient to support the development and maintenance of complex software systems. It is necessary to design a new methodology to better prepare and organize the important knowledge missed by LLMs.

### v2

@@ 1. Introduction
Recent advancements in large language models (LLMs) have demonstrated their impressive capabilities in software engineering (SE) tasks, such as code completion, bug detection, and software documentation. These models excel at understanding and generating code, streamlining development cycles, and assisting developers with routine tasks. Leveraging vast amounts of public domain data, LLMs have notably improved productivity and code quality in simpler, well-defined software systems. However, amidst rapid AI advancements dominated by LLMs, claims about “generative AI ending programming” and “revolutionizing software paradigms” abound. This hype may mislead enterprise leaders into expecting “10x efficiency gains”, leading to unrealistic demands on development teams. Consequently, investments flood into LLM fine-tuning and tool procurement, while essential long-term initiatives, such as building robust CI/CD pipelines, digital foundations, and knowledge repositories, are neglected. Current research (including industry efforts) excessively prioritizes the “glamour effect” of LLMs and AI agents, paying scant attention to software engineering’s intrinsic challenges.
In enterprise software development, tasks predominantly revolve around maintenance. Even “adding new code” is essentially a maintenance task: new code is rarely written on a blank canvas; instead, it must be integrated into proper positions within an existing software architecture through targeted additions and modifications. This process demands rigorous change impact analysis and regression testing to mitigate potential side effects. As software systems increase in complexity, LLMs face significant challenges. While they are effective in tackling individual tasks, their ability to support the maintenance and evolution of large, intricate systems is limited. LLMs struggle to capture the tacit knowledge essential to understanding complex software systems—such as design rationales, high-level functionalities, and system interdependencies—which is often undocumented or informally communicated. This challenge is exacerbated by knowledge decay, where critical context gradually fades or becomes fragmented across developer memories, scattered discussions, and incomplete documentation. As a result, LLMs lack the holistic, system-wide perspective needed to manage the intricate relationships and ongoing changes inherent in software evolution.
[…]
To address these challenges, we propose the concept and framework of Code Digital Twin, a conceptual representation of tacit knowledge that encapsulates the concepts, functionalities, and design rationales behind code elements, co-evolving with the software. Code Digital Twin addresses knowledge decay by extracting, accumulating, and continuously updating high-level knowledge about code through bidirectional mapping and co-evolution with the codebase. The framework structures tacit knowledge into two key categories: the artifact-oriented backbone, which establishes a foundational structure by linking concepts and functionalities to software artifacts, and the rationale-centric explanations, which enrich this foundation by capturing the decisions that have shaped the system’s design and evolution. We elevate software knowledge accumulation to a higher purpose: codifying knowledge about such ultra-complex systems becomes part of humanity’s civilizational record.
@@
The methodology for realizing a Code Digital Twin combines structured and unstructured knowledge representations, integrating knowledge graphs, frames, and textual descriptions to capture domain concepts, functionalities, and rationales. The construction pipeline consists of three key phases: artifact-oriented backbone knowledge extraction, which employs top-down schema-based and schema-free extraction from documentation and bottom-up functionality summarization from source code; rationale-centric explanatory knowledge extraction, where LLMs extract and structure design rationales from unstructured sources like mailing lists and discussions using predefined knowledge frame schemas; and artifact-knowledge reflection construction, which links extracted knowledge across software artifacts and elements to create a cohesive representation. This pipeline supports incremental updates, ensuring that the Code Digital Twin evolves in sync with software changes, maintaining continuous alignment with the latest development state.
@@
By providing integrated knowledge and ensuring continuous alignment with the evolving software, the Code Digital Twin enhances collaboration between LLMs and developers. It empowers LLMs to assist in software maintenance through various modes, including software question answering, context-aware development co-pilot, and autonomous development auto-pilot. This framework presents an innovative approach for the systematic accumulation of tacit knowledge in long-term software evolution, enabling more efficient maintenance and adaptation over time.
@@ 2. Motivation
Software systems are growing increasingly complex, while the pool of developers continues to evolve and fluctuate. This dynamic landscape makes software development more challenging, with essential knowledge becoming harder to retain and apply. A notable example is the Linux kernel, which remains under rapid development (Linux, 2025), with over 243,000 contributors producing more than 1,060,172 commits in 2021 alone. Its global, collaborative ecosystem introduces significant complexity, compounded by the continual influx and departure of developers. Software maintenance, a vital yet difficult phase of the development lifecycle, involves tasks such as diagnosing issues, assessing their impact, and integrating new code into existing architectures. These tasks require a deep understanding of both the system’s internal logic and external operating contexts. Despite recent advancements, evaluations such as those from OpenAI (Samuel et al., 2025) indicate that state-of-the-art LLMs remain limited in their ability to address complex, real-world maintenance challenges effectively.
2.1. Challenges in Complex Maintenance Tasks
Tacit Knowledge is Difficult to Capture. A significant portion of essential software knowledge is tacit—residing not in documentation or code, but in the minds of developers or in informal artifacts like whiteboards and meeting discussions. While LLMs excel at leveraging publicly available corpora, they falter when such tacit knowledge is missing or poorly documented. Critical insights into design rationale, historical decisions, and architectural intent are often unavailable for machine learning systems to absorb. This knowledge is also domain-specific, continuously evolving, and often requires human interpretation—making it unsuitable for passive ingestion via fine-tuning alone.
Internal and External Complexity. Software complexity manifests in both internal structures and external dependencies. Internally, systems are built upon intricate design decisions that result in deep interdependencies across the codebase. Externally, they must operate within a broader ecosystem—hardware, networks, runtime environments, and user scenarios—all of which influence design and maintenance. Contrary to simplified metaphors like “pattern soup”, software is less like a collection of reusable motifs and more akin to a precision instrument: highly engineered, tightly coupled, and often opaque. Its complexity exceeds that of mechanical systems due to four defining traits:
  • Inherent Ambiguity: Software encodes logic and intent, where small errors can have cascading effects, and boundaries between components are often ill-defined.
  • Contextual Dependencies: Code interacts with systems and conditions beyond itself, making isolated reasoning insufficient.
  • Internal Entanglement: High-level requirements or bug reports may map ambiguously to deeply entangled code, complicating traceability.
  • Evolutionary Uniqueness: Software evolves through decentralized contributions, leading to technical debt, duplication, and legacy structures.
LLMs lack the high-level, abstract reasoning and system-level perspective necessary to navigate such intricacies. They are not equipped to internalize the broader context in which systems operate or the creative processes that drive architectural decisions.
The “LLM Catch-All” Fallacy. A prevalent misconception in the field is the belief that LLMs can serve as a universal solution for software engineering by simply ingesting all relevant knowledge through pretraining or fine-tuning. This so-called LLM catch-all view treats software as a sequence of patterns, suggesting that intelligent development is merely a matter of recognizing and reproducing them. It implies that “all code is derivative”, downplaying the originality, structure, and judgment involved in real-world software work. However, this approach has fundamental flaws:
  • Flattened Information: Code without contextual metadata lacks essential abstractions and decision-making traces.
  • Absence of Design Intelligence: LLMs cannot infer the creative, system-wide reasoning behind design decisions from isolated tokens.
To enable intelligent support for tasks like root cause analysis, change planning, or impact estimation, we must move beyond superficial pattern recognition. These tasks require System 2 thinking—deliberate, structured, and cognitively intensive reasoning—something current LLMs are not equipped to perform independently.
2.2. Toward a New Methodology: Code Digital Twin
Given these challenges, LLMs are insufficient to support the development and maintenance of complex software systems on their own. Hence, the LLM catch-all approach fails for complex software. A new methodology is needed to systematically capture, organize, and maintain the tacit knowledge that underpins software evolution. The Code Digital Twin framework addresses this need by providing a structured representation of software knowledge that co-evolves with the system. Success hinges on:
  • Continuously codifying knowledge from development and evolution.
  • Structuring knowledge (e.g., via graphs).
  • Enabling visible, collaborative curation by developers.
Code Digital Twin is knowledge engineering for ultra-complex systems. But why “digital twin” for already-digital software? Code is the physical layer (files, classes, functions), while the Digital Twin captures the logical layer (requirements, design, decisions)—mapping intent to implementation. For cloud-native systems, we extend this to System Digital Twin, encompassing deployment topology, runtime environments, configuration states, and live updates. For monolithic systems (e.g., Linux), Code Digital Twin suffices—focusing on high-level abstractions and evolution. By continuously codifying knowledge, structuring it via graphs and frames, and enabling collaborative curation by developers, the Code Digital Twin empowers LLMs to assist in complex maintenance tasks more effectively, ensuring that critical knowledge is preserved and accessible for the long-term sustainability of software systems.